OpenStudy (anonymous):

In a survey, 36% of the students said thee keep an electronic diary. There were 900 students who kept an electronic diary. How many students were in the survey?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

900/36 *100 =25*100=2500

OpenStudy (anonymous):

R u sure?

OpenStudy (aripotta):

yea, that's right

OpenStudy (anonymous):

see out of 100% only 36% have that electronic diary so 900 is just 36% and we need to find 100%

OpenStudy (aripotta):